OpenStudy (silverlark2):

In curve y= 2(x + a)(x + b) cuts the x-axis at points A and B where a and b are constants. Given that the coordinates of A and B are (-6,0) and (2,0) respectively, find the equation of the quadratic graph.

OpenStudy (saifoo.khan):

Make 2 equations.. and insert the respective set values given

OpenStudy (saifoo.khan):

y= 2(x + a)(x + b) 0 = 2(-6 + a)(-6 + b) same as with second one

OpenStudy (silverlark2):

Thank you!

OpenStudy (saifoo.khan):

Np.

OpenStudy (silverlark2):

Hi sorry I still don't know how to carry on from there, do I expand 0=(2-6 + a)(-6 + b) and form a simultaneous equation? Very lost >< Thanks